OpenStudy (anonymous):

Find the value of x that makes the equation true. –6x = 42 A. –7 B. –252 C. 48 D. 36

OpenStudy (anonymous):

On questions like these, if you are unsure as to how to solve for x, it can be helpful to simply substitute answers in for the x value and see if they make the statement true.

OpenStudy (anonymous):

I think its a because if you substitute x with 7 ur answer will be 42

OpenStudy (anonymous):

so for instance: if we substitute in answer a. for x you get something like this: -6 (-7) = 42 and then solve, 42=42, this is true so a is the correct answer.

OpenStudy (anonymous):

Identify which property of equality is used to transform Equation 1 to Equation 2. Equation 1: 3x = 12 Equation 2: x = 4 A. subtraction property of equality B. division property of equality C. addition property of equality D. multiplication property of equality

OpenStudy (anonymous):

however if you are not given answers, it is important to know how to solve this equation algebraically: -6x=42 the equation you start with -6x/6 = 42/-6 this is the division property of equality (what you do to one side, you do to the other.) and x=-7

OpenStudy (anonymous):

D.because if u replace the x with 4 your answer will be d multiplication if u multiply 3 times 4 wich will give u 12

OpenStudy (anonymous):

you have to divide by x to find the answer so the b division property of equality is the correct answer

OpenStudy (anonymous):

the division property of equality is used to find the answer

OpenStudy (anonymous):

you can divide that will be easier 4 didvide by 12 will be 3 but wont you find the same anwer if u multiply. 3 times 4=12

OpenStudy (anonymous):

you will get the same answer, 3(4)=12 and 12/3=4 and 12/4=3. In order to change 3x=12 into x=4 with out knowing the answer simple math, you must divide, this is called the division property of equality

OpenStudy (anonymous):

Ok thx

OpenStudy (anonymous):

yeah, let me know if you have any other questions!